OpenStudy (anonymous):

ANSWER IN 5 MINUTES FOR A MEDAL AND FAN! What is the y-intercept of the line? y = 6x - 4

OpenStudy (igreen):

The equation is in the form of y = mx + b. Where m = slope, and b = y-intercept. So what's the y-intercept? @im.celibate

OpenStudy (anonymous):

2.

OpenStudy (anonymous):

Because 6 - 4 = 2

OpenStudy (igreen):

Where did you get 2 from? y = mx + b ^ ^ y-intercept | Slope y = 6x - 4

OpenStudy (igreen):

No, you don't do anything, the equation tells you what the answer is.

OpenStudy (anonymous):

i got it from my mom. Sorry XD

OpenStudy (igreen):

Lol, that's okay.

OpenStudy (igreen):

y = mx + b M = Slope B = y-intercept y = 6x + b ? = Slope ? = y-intercept?

OpenStudy (anonymous):

I still don't know. Lol. I am horrible at math.

OpenStudy (anonymous):

So the y-intercept would just be y = mx + b

OpenStudy (igreen):

Since b is the y-intercept in y = mx + b, the last number in y = 6x - 4 is going to be the y-intercept, which is -4.